77% of adults cannot name a supreme court justice. you randomly select five adults and ask each if they can name any of the justices. the random variable represents the number of adults who cannot name a supreme justice. find the mean of the binomial distribution
Answer #1

solution:

given

p=0.77

n=5

the mean of the binomial distribution=E(X)=n*p

=5*0.77

=3.85